You want to purchase computer system for your newly established business. The system you desire to purchase is priced at Rs. 50,000. At store sales person offers you that instead of buying at full price today, you can divide your payment over 2 years. If you accept the installment plan, you will have to pay Rs. 50,000 ÷ 24 per month plus an administrative fee of Rs.40 per month to be added to your monthly bill.

If the on-going interest rate is 18%, which option do you thinks is better for you?



1
Expert's answer
2021-04-30T10:48:45-0400

Future value of 50,000 payed today is 50,000×(1+0.18)^2=69,620.

If to divide 50,000 over 2 years, you should pay 50,000/24+40=2123.33 per month. Interest rate per month is 0.18/12=0.015. Future value of the installment plan is 2123.33×(1+0.015)^24+2123.33×1.015^23+.....+2123.33×1.015=61,710. So, this plan is better